OpenStudy (anonymous):

Nicole has a clay vase shaped as a triangular prism with a base area of 136.9 square inches. It is one third filled with water. If there are 575 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth? 13.5 inches 8.3 inches 12.6 inches 7.1 inches

OpenStudy (anonymous):

(1/3) b * h = 575 => h = 3* 575 / 136.9 = 12.6 in Hope this helps!

OpenStudy (anonymous):

it does thank you! :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

@Wellplayedsir  This is a " TRIANGULAR PRISM" "(1/3) b * h = V" is used for volume of PYRAMIDS I think you wrongly typed the question !

OpenStudy (anonymous):

What do you think @ZhangYan ?

OpenStudy (anonymous):

I didnt write the question wrong thats how it was given to me on my worksheet, but thank you for your input